Consider two neighboring island countries called Dolorium and Arcadia. They each have 4 million labor hours available per week that they can use to produce jeans, corn, or a combination of both. The following table shows the amount of jeans or corn that c
Answer #1

Blanks:-

1) 20/5 =4 pairs

2) 16/8 =2 pairs

3) Contente

4) Dolorium

5) 8*4 =32

6) 20*4 = 80

7) 32-23 = 9

8) 80-68= 12

18 million corn is being exchanged for 54 million jeans

Dolorium

Contente

Corn

Jeans

Corn

Jeans

Without Trade

Production

15

20 8 48

Consumption

15

20 8 48

With Trade

Production

0

80 32 0

Trade action

Import 18

Export 54 Export 18 Import 54

Consumption

18

80-54= 26 32-18 = 14 54

Gains from Trade

Increase in Consumption

18-15 =3

26-20 =6 14-8 = 6 54-48 = 6